OpenStudy (amyna):

y=-3(sec^2x - tan^2(x)) + 4x^3 I do not know how to solve this! Please help with the calc problem!! Thank you!

OpenStudy (dan815):

sec(x) = 1/cos(x)

OpenStudy (freckles):

recall Pythagorean identities \[\sin^2(x)+\cos^2(x)=1 \\ \text{ divide both sides by } \cos^2(x) \\ \tan^2(x)+1=\sec^2(x) \\\] anyways I don't know what it is meant by solve

OpenStudy (freckles):

could you say what solve means here @amyna if you are here

OpenStudy (freckles):

you say this is a cal problem do you really mean differentiate are you trying to find the inverse like what are you doing exactly

OpenStudy (amyna):

yup sorry! differentiate it, thanks!

OpenStudy (freckles):

so you need help to differentiate -3+4x^3 just use constant rule and constant multiple rule and power rule?

OpenStudy (freckles):

-3 is a constant the derivative of a constant is 0

OpenStudy (freckles):

\[\frac{d}{dx} cx^n \\ c \frac{d}{dx} x^n \text{ by constant multiple rule } \\ c n x^{n-1} \text{ by power rule }\]

OpenStudy (freckles):

\[\frac{d}{dx}(-3)=? \\ \frac{d}{dx}4x^3=?\]

OpenStudy (amyna):